1. Find the value of the finite population correction factor when:

 

a)   n = 10 and N = 10000

b)   n = 40 and N= 10,000


1
Expert's answer
2022-04-07T08:32:58-0400

a) "FPC=\sqrt{\cfrac{N-n}{N-1}}=\sqrt{\cfrac{10000-10}{10000-1}}=\sqrt{\cfrac{9990}{9999}}=0.99955."


b) "FPC=\sqrt{\cfrac{N-n}{N-1}}=\sqrt{\cfrac{10000-40}{10000-1}}=\sqrt{\cfrac{9960}{9999}}=0.99805."